# FedAPTA: Federated Multi-task Learning in Computing Power Networks with Adaptive Layer-wise Pruning and Task-aware Aggregation

### v1

@@ III-D Effects of Model Similarity Metrics
TABLE VI: Existing work in FL.
@@
Category Frameworks Main Techniques Under CPNs Multi-task Model Pruning
FedACT [6] Adaptive chained training $\checkmark$ $\times$ $\times$
FL in CPNs Lin et al. [7] Incentive mechanism based on Stackelberg game $\checkmark$ $\times$ $\times$
FedTAR [4] Dynamic network structure and collaboration frequency $\checkmark$ $\times$ $\times$
@@
FedDrop [9] Randomly drop to obtain a sub-model $\times$ $\times$ $\checkmark$
@@
FedRolex [10] Uses a rolling window to extract parts of a global model $\times$ $\times$ $\checkmark$
@@
FL with pruning Hermes [11] Applies structured pruning to obtain a sub-model $\times$ $\times$ $\checkmark$
@@
FedMP [12] Dynamic pruning ratios via multi-armed bandit algorithm $\times$ $\times$ $\checkmark$
FedLPS [13] Adaptive channel-wise pruning $\times$ $\checkmark$ $\checkmark$
@@
FedEM [38] Hybrid distribution modeling $\times$ $\checkmark$ $\times$
@@
MTFL [39] Allows devices to train personalized models $\times$ $\checkmark$ $\times$
@@
FL in multi-task FedMSplit [40] Dynamic graphs for relationships in multimodal models $\times$ $\checkmark$ $\times$
@@
scenarios MOCHA [41] Modeling inter-device relationships $\times$ $\checkmark$ $\times$
@@
Ghosh et al. [42] Group devices by data distribution through robust clustering $\times$ $\checkmark$ $\times$
@@
CFL [15] Device clustering based on gradients $\times$ $\checkmark$ $\times$
@@
ClusterFL [43] Uses a cluster indicator matrix Selects relevant devices $\times$ $\checkmark$ $\times$
FedDGA [44] Dynamic guided attention $\times$ $\checkmark$ $\times$
IV Related Work
IV-A Federated Learning in Computing Power Networks
In CPNs, the distributed architecture and privacy-preserving feature of FL present inherent advantages, making it highly suitable for processing data on widely distributed devices. FL is a distributed ML framework that is designed to enable the collaborative training of a global model across numerous devices. Importantly, during the FL process, only the models are transmitted to the central server, without the need to transmit any raw local data [1], thereby ensuring data privacy protection.
In recent years, many related studies have been proposed to implement FL under CPNs. For example, FedACT [6] improves the efficiency and accuracy of FL in CPNs through an adaptive chained training approach. Lin et al. [7] proposed an incentive mechanism to address the issue of insufficient device participation motivation. FedTAR [4] dynamically adjusts each device’s model structure and communication frequency based on task complexity and computing resources, significantly reducing overall system energy consumption while ensuring model accuracy. Although these studies improve FL from different aspects, they overlook the aspect of multi-task learning in CPNs.
IV-B Model Pruning in Federated Learning
@@
With the continuous advancement of deep neural networks (DNNs), the increasing number of parameters and the resulting computational overhead have significantly slowed down model training. To tackle this pressing challenge, a pruning method has been introduced by [45], aiming to compress DNNs by removing redundant parameters and structural components. This approach seeks to reduce both computational and storage costs, while maintaining model performance [16, 46, 47, 48].
@@
In FL, the large size of models has emerged as a significant bottleneck, impeding their efficient deployment and training on resource-constrained devices [12]. Since the central server cannot directly access local data, traditional pruning strategies [45, 49] that depend on global data distribution are difficult to apply directly in FL settings. To address this limitation, prior research has investigated incorporating model pruning into FL frameworks as a means to reduce the number of model parameters. For example, FedDrop [9] trains only a subset of the global model on each device to lower local computation costs. FedRolex [10] utilizes a sliding window mechanism to extract sub-models, enabling balanced training across different parts of the global model; Hermes [11] applies structured pruning on devices to obtain lightweight sub-models. FedMP [12] employs a multi-armed bandit algorithm to dynamically adjust the pruning ratio per device. FedLPS [13] introduces an adaptive channel-wise pruning algorithm to produce task-specific lightweight predictors. However, most existing approaches adopt a uniform pruning ratio across all candidate layers, without considering the varying importance of individual layers within the model.
IV-C Federated Multi-Task Learning
@@
In practical FL applications, it is common for devices to perform different tasks. Existing studies, such as Hermes [11], have shown that in such multi-task learning scenarios, the performance of a globally trained model may fall short of that achieved by models trained locally, which fundamentally contradicts the core objective of FL.
@@
To address this issue, several works have explored the development of personalized models. For instance, FedEM [38] models each device’s data distribution as a mixture of multiple latent components, enabling the joint learning of shared component models and personalized mixture weights, which significantly enhances both model accuracy and fairness. MTFL [39] introduces non-federated batch normalization layers, allowing devices to train personalized deep neural networks, thereby improving accuracy and convergence speed. FedMSplit [40] employs a dynamic graph structure to represent relationships among multimodal device models, enabling the learning of personalized yet globally correlated local models. However, it is important to note that these approaches primarily focus on training models tailored to individual devices.
@@
A few studies have further explored multi-task learning in the context of FL. For example, MOCHA [41] applies a multi-task learning framework to train multiple models across devices. Ghosh et al.[42] propose a robust clustering algorithm that groups devices with similar data distributions while resisting interference from Byzantine devices. CFL[15] clusters devices based on the similarity of their gradient updates, enabling the learning of specialized models within each cluster—particularly beneficial in non-i.i.d. settings. ClusterFL [43] enhances training efficiency by discarding straggling devices and selecting those with greater relevance through a cluster indicator matrix. FedDGA [44] addresses non-i.i.d. data distribution and task imbalance issues by introducing Dynamic Guided Attention (DGA) for adaptive feature alignment and Dynamic Batch Weighting (DBW) for loss balancing, achieving a high accuracy while maintaining communication efficiency.
V Conclusion
@@
In this paper, we propose FedAPTA, a FL framework designed to address the issues of computing resources wastage and suboptimal global model performance caused by directly aggregating local models from different tasks in CPNs. To mitigate the first issue, we design an adaptive layer-wise model pruning method in FedAPTA, which permits heterogeneous devices to participate in federated training with pruned models. Furthermore, FedAPTA introduces a novel heterogeneous model recovery algorithm, which is capable of effectively aggregating pruned models with the assistance of the latest global model information. To alleviate the second issue, FedAPTA achieves task-aware model aggregation by performing device clustering. Specifically, it enables local models to be aggregated separately for each task. Extensive experimental comparisons demonstrate that FedAPTA is capable of preserving superior model accuracy across a range of FL scenarios while avoiding unnecessary computing resource consumption. In this future, we plan to explore FedAPTA’s performance in other task domains beyond classification.

### v2

@@ 3.4 Effects of Model Similarity Metrics
(a) Non-i.i.d.
(b) i.i.d.
Table 6: Existing work in FL.
@@
Category Frameworks Main Techniques Multi-task Model Pruning
@@
FedDrop [caldas2018expanding_feddrop] Randomly drop to obtain a sub-model $\times$ $\checkmark$
@@
FedRolex [alam2022fedrolex] Uses a rolling window to extract parts of a global model $\times$ $\checkmark$
@@
FL with pruning Hermes [li2021hermes] Applies structured pruning to obtain a sub-model $\times$ $\checkmark$
@@
FedMP [jiang2022fedmp] Dynamic pruning ratios via multi-armed bandit algorithm $\times$ $\checkmark$
FedLPS [jia2024fedlps] Adaptive channel-wise pruning $\checkmark$ $\checkmark$
@@
FedEM [marfoq2021federated] Hybrid distribution modeling $\checkmark$ $\times$
@@
MTFL [mills2021multi] Allows devices to train personalized models $\checkmark$ $\times$
@@
FL in multi-task FedMSplit [chen2022fedmsplit] Dynamic graphs for relationships in multimodal models $\checkmark$ $\times$
@@
scenarios MOCHA [smith2017federated_mocha] Modeling inter-device relationships $\checkmark$ $\times$
@@
Ghosh et al. [ghosh2019robust_gho] Group devices by data distribution through robust clustering $\checkmark$ $\times$
@@
CFL [sattler2020clustered_cfl] Device clustering based on gradients $\checkmark$ $\times$
@@
ClusterFL [ouyang2021clusterfl] Uses a cluster indicator matrix Selects relevant devices $\checkmark$ $\times$
FedDGA [10382491] Dynamic guided attention $\checkmark$ $\times$
4 Related Work
4.1 Federated Learning
FL is a distributed ML framework that is explicitly designed to enable the collaborative training of a global model across multiple devices, while simultaneously preserving and safeguarding data privacy [mcmahan2017communicationfedavg]. Within the paradigm of FL, each participating device utilizes its locally stored data to perform a number of local training steps, after that, the updated model parameters are transmitted to a central server. The central server subsequently performs an aggregation operation over all the received parameters from the various devices, thereby updating the shared global model. Importantly, throughout this entire process, only the updated models are communicated to the central server, and there is no need to transmit any raw local data.
However, when deploying deep neural networks (DNNs) in the FL framework, the system often incurs substantial communication and computation overhead [huang2017densely_dnn]. Various efforts have attempted to alleviate this issue through a number of techniques, such as reducing the frequency of model aggregation on the central server [mcmahan2017communicationfedavg], applying parameter sparsification [han2020adaptive_dense1, hu2021federated_dense2_new], and quantization [sattler2019robust_liang1, hubara2018quantized_liang2] techniques. However, these approaches have predominantly focused on improving communication efficiency, while overlooking the fact that devices often exhibit limited and heterogeneous computational capabilities [gao2022survey_hetero]. This heterogeneity among devices typically leads to performance degradation in real-world FL. This paper focuses on the impact of this computational capability heterogeneity on FL.
4.2 Model Pruning in Federated Learning
@@
With the continuous advancement of DNNs, the increasing number of parameters and the resulting computational overhead have significantly slowed down model training. To tackle this pressing challenge, a pruning method has been introduced by [han2015learning_pruningst], aiming to compress DNNs by removing redundant parameters and structural components. This approach seeks to reduce both computational and storage costs, while maintaining model performance [liu2017learning_pr1, he2017channel_pr2, han2015learning_prCompressParams, lin2018accelerating_prStr].
@@
In FL, the large size of models has emerged as a significant bottleneck, impeding their efficient deployment and training on resource-constrained devices [jiang2022fedmp]. Since the central server cannot directly access local data, traditional pruning strategies [han2015learning_pruningst, fang2023depgraph_pruning] that depend on global data distribution are difficult to apply directly in FL settings. To address this limitation, prior research has investigated incorporating model pruning into FL frameworks as a means to reduce the number of model parameters. For example, FedDrop [caldas2018expanding_feddrop] trains only a subset of the global model on each device to lower local computation costs. FedRolex [alam2022fedrolex] utilizes a sliding window mechanism to extract sub-models, enabling balanced training across different parts of the global model; Hermes [li2021hermes] applies structured pruning on devices to obtain lightweight sub-models. FedMP [jiang2022fedmp] employs a multi-armed bandit algorithm to dynamically adjust the pruning ratio per device. FedLPS [jia2024fedlps] introduces an adaptive channel-wise pruning algorithm to produce task-specific lightweight predictors. However, most existing approaches adopt a uniform pruning ratio across all candidate layers, without considering the varying importance of individual layers within the model.
4.3 Federated Multi-Task Learning
@@
In practical FL applications, it is common for devices to perform different tasks. Existing studies, such as Hermes [li2021hermes], have shown that in such multi-task learning scenarios, the performance of a globally trained model may fall short of that achieved by models trained locally, which fundamentally contradicts the core objective of FL.
@@
To address this issue, several works have explored the development of personalized models. For instance, FedEM [marfoq2021federated] models each device’s data distribution as a mixture of multiple latent components, enabling the joint learning of shared component models and personalized mixture weights, which significantly enhances both model accuracy and fairness. MTFL [mills2021multi] introduces non-federated batch normalization layers, allowing devices to train personalized deep neural networks, thereby improving accuracy and convergence speed. FedMSplit [chen2022fedmsplit] employs a dynamic graph structure to represent relationships among multimodal device models, enabling the learning of personalized yet globally correlated local models. However, it is important to note that these approaches primarily focus on training models tailored to individual devices.
@@
A few studies have further explored multi-task learning in the context of FL. For example, MOCHA [smith2017federated_mocha] applies a multi-task learning framework to train multiple models across devices. Ghosh et al.[ghosh2019robust_gho] propose a robust clustering algorithm that groups devices with similar data distributions while resisting interference from Byzantine devices. CFL[sattler2020clustered_cfl] clusters devices based on the similarity of their gradient updates, enabling the learning of specialized models within each cluster—particularly beneficial in non-i.i.d. settings. ClusterFL [ouyang2021clusterfl] enhances training efficiency by discarding straggling devices and selecting those with greater relevance through a cluster indicator matrix. FedDGA [10382491] addresses non-i.i.d. data distribution and task imbalance issues by introducing Dynamic Guided Attention (DGA) for adaptive feature alignment and Dynamic Batch Weighting (DBW) for loss balancing, achieving a high accuracy while maintaining communication efficiency.
5 Conclusion
@@
In this paper, we propose FedAPTA, a FL framework designed to address the challenge of heterogeneous computational capabilities and the suboptimal global model performance caused by directly aggregating local models trained on different tasks. To mitigate the first issue, we design an adaptive layer-wise model pruning method in FedAPTA, which permits heterogeneous devices to participate in federated training with pruned models. Furthermore, FedAPTA introduces a novel heterogeneous model recovery algorithm, which is capable of effectively aggregating pruned models with the assistance of the latest global model information. To alleviate the second issue, FedAPTA achieves task-aware model aggregation by performing device clustering. Specifically, it enables local models to be aggregated separately for each task. Extensive experimental comparisons demonstrate that FedAPTA is capable of preserving superior model accuracy across a range of FL scenarios. In this future, we plan to explore FedAPTA’s performance in other task domains beyond classification.